OpenStudy (anonymous):

A stone is thrown upward from ground level. With what minimum speed should the stone be thrown so as to reach a height of 64 feet?

OpenStudy (anonymous):

x = x0 + v0t + a/2 t^2. 64 = 0 + v0t -9.8/2 t^2. (1) v0 is what you want. You must solve for time first. V = v0 + at. 0 = v0 -9.8t . (2) Therefore v0 = 9.8 t. Put this into (1) to find t and then plug that time into either equation to find v0.

OpenStudy (anonymous):

So 64 = 9.8t^2 - 4.9t^2 .

OpenStudy (anonymous):

9.8t is for meters, but i'll change the formula for feet and solve it myself. thank you!

OpenStudy (anonymous):

Sure thing.